# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

### v1

@@ A.1.6 Grid stability and management cost savings (GSMS)
Reduced DSO/TSO costs for grid stability arise from several ODP-enabled improvements (Section 3.1). We calculate the annual benefit as
where is the ODP-enabled forecasting improvement fraction (informed by smart-grid pilot results [53, 54])888We assume , , , , , in our calculations., is the total number of EVs, is the total number of ETs, [MWh/yr] and [MWh/year] are the annual charging energy per EV and per ET, respectively, and [€/MWh] is the unit fuel-saving value. The product captures savings from reduced fuel use due to improved timing of EV/ET charging. The term represents avoided operational-downtime costs: is the baseline downtime cost, [MW] is installed RES capacity, and is the RES full-load hours, so that multiplying by estimates the portion of downtime cost deferral attributable to RES-driven stability improvements.
Finally, is the ODP-enabled grid-stability savings fraction. Multiplying by where [€/MWh] is the market-price reduction rate, yields additional benefits from reduced imbalance and congestion charges when EV/ET charging shifts under ODP’s real-time control.
@@ A.1.7 CO2 emission reduction benefits
Monetized GHG emission reductions arise from ODP’s AI-driven increase in RES utilization and optimized EV/ET charging (Section 3.1). We model the annual benefit as
where is the ODP-enabled reduction fraction999we take 12%, 0.4 MWh/year, 0.8 MWh/year,0.025 tonne CO2/MWh, [€/tonne CO2]=50 €/tonne, in our calculation., is the total number of EVs and [MWh/year per vehicle] is the annual energy equivalent per EV after AI-optimized charging; similarly, is the total number of ETs and [MWh/year per vehicle] is the annual energy equivalent per ET. The factor [tonne CO2/MWh] is the emission factor, and [€/tonne CO2] is the CO2 price following EU guidance [50]. Thus, gives the combined annual MWh associated with RES, EV, and ET operation; multiplying by yields total tonnes of CO2 avoided, and converts to euros, with capturing the the incremental reduction enabled by ODP.
@@ A.1.8 Reduction in air pollutant (RAP)
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and transport (Section 3.1). We model the annual benefit as
where [kg/km] is the EV-air pollutant emission factor101010We take 0.000005 kg/km, 20,000 km/year, 20%, 1%, 10%, 40,000 km/year, 12%, 0.025 kg/MWh in our calculations., [km/yr] is the annual distance per EV, and is the total number of EVs, so that gives the total annual EV-related pollutant reduction in kilograms. Here, is the fraction of RES capacity initially considered each year for the emitting pollutants calculation, is the fraction of RES output shifted through DR, and is the grid decarbonization factor; therefore represents the annual MWh of RES energy shifted and decarbonized. Similarly, [kg/km] is the ET-air pollutant emission factor, [km/year] is the annual distance per ET, is the total number of ETs, and is the ODP-enabled emission reduction fraction, so that gives the ET-related pollutant reduction in kilograms. The term [kg/MWh] is the average emission factor for SO2, NO2, and PM10, making the annual kilograms of pollutants avoided by cleaner RES generation. Finally, [€/kg] is the unit cost of local air pollutants based on EU guidelines [56].
@@ A.3 OPEX breakdown
This appendix details the OPEX components for the AI-driven ODP. The figures in Table 10 represent the discounted present value (PV) of recurring annual costs over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The Annual average PV provides an annualized perspective of these costs, calculated from the total 10-year discounted OPEX. Costs specifically related to sustaining and evolving the AI engine are explicitly considered.
@@
A.4 Per-country CBA summary and Monte Carlo simulation (MCS) results
This appendix provides a summary of the final CBA outcomes and Monte Carlo Simulation (MCS) results for the AI-driven ODP, disaggregated for each of the three case study countries: Austria (AT), Hungary (HU), and Slovenia (SI). The numerical values presented here are the actual calculated results based on country-specific assumptions and scaled costs, demonstrating the national-level impacts derived from the detailed financial and operational models.
| Metric | Austria (AT) | Hungary (HU) | Slovenia (SI) |
| Deterministic results (Base Case) | |||
| Total benefits (PV) | 550.8 | 434.7 | 248.4 |
| Total costs (PV) | 390.4 | 310.8 | 176.0 |
| NPV (PV) | 160.4 | 123.9 | 72.4 |
| BCR | 1.41 | 1.40 | 1.41 |
| Monte Carlo simulation (Mean values from 10,000 Trials) | |||
| Mean NPV (PV) | 168.2 | 127.5 | 74.3 |
| Mean BCR | 1.43 | 1.42 | 1.42 |
| Dominant benefit categories (% of national total benefits) | |||
| Economic (ROD, AEC, FES, ROETAS, CSDR-PLR) | |||
| Environmental (CO2, RAP) | |||
| Reliability (GSMS) | |||
1.
Note: These numerical values are derived from detailed country-specific financial and operational models, reflecting the specific growth rates, market prices, and grid characteristics outlined in Table 5 and other appendices. The slight variations in BCR across countries reflect the unique cost structures and benefit profiles of each national context. Dominant benefit categories highlight the primary drivers of value in each country, influenced by their respective energy and transport system characteristics.
@@ B.3 Annual costs of the ODP (aggregated)
Note: Annual CAPEX and OPEX are discounted sums over the 2026–2035 period, aggregated across Austria, Hungary, and Slovenia. The Total row shows the ten-year aggregate costs for each category. In addition, a one-time upfront fixed cost of €89.6 million is incurred in 2026 to cover connectivity (5G/LPWAN), communication protocols and message queues, databases and data lakes, data integration middleware, market integration middleware, AI models for grid optimization, demand response tools, route optimization engines, weather-forecasting algorithms, software deployment and integration.

### v2

@@ A.1.6 Grid stability and management cost savings (GSMS)
Reduced DSO/TSO costs for grid stability arise from several ODP‐enabled improvements (Section 3.1). Grid congestion and reliability benefits are quantified through reductions in expected energy not served and congestion events. Baseline reliability is estimated from national outage statistics and grid development reports for Austria, Hungary, and Slovenia Statista Research Department [2023]. The ODP’s impact is modeled as the incremental reduction in expected energy not served, reflecting improved flexibility and RES integration. Monetization is based on country-specific Values of Lost Load, as reported in regulatory and market studies. Thus, we calculate the annual benefit of GSMS as
where is the ODP‐enabled forecasting improvement fraction (informed by smart‐grid pilot results [U.S. Department of Energy, Office of Electricity, 2019, Electric Power Research Institute (2020), EPRI]), is the total number of EVs, is the total number of ETs, [MWh/yr] and [MWh/year] are the annual charging energy per EV and per ET, respectively, and [€/MWh] is the unit fuel‐saving value. The product captures savings from reduced fuel use due to improved timing of EV/ET charging. The term represents avoided operational‐downtime costs: is the baseline downtime cost, [MW] is installed RES capacity, and is the RES full‐load hours, so that multiplying by estimates the portion of downtime cost deferral attributable to RES‐driven stability improvements.
Finally, is the ODP‐enabled grid‐stability savings fraction. Multiplying by where [€/MWh] is the market‐price reduction rate, yields additional benefits from reduced imbalance and congestion charges when EV/ET charging shifts under ODP’s real‐time control.
@@ A.1.7 CO2 emission reduction benefits
Monetized GHG emission reductions arise from ODP’s AI‐driven increase in RES utilization and optimized EV/ET charging. To increase granularity, we now differentiate both temporally and spatially. We model the annual benefit as
where is the ODP‐enabled reduction fraction, [European Commission, 2022]. where is the ODP‐enabled reduction fraction varying by time and country, is the number of EVs in operation in country at time and [MWh/vehicle] is their average energy demand after AI‐optimized charging. Similarly, is the number of ETs, and [MWh/vehicle] is their annualized energy demand. denotes the renewable generation (capacity multiplied by capacity factor) available in country at time . The factor [tonne CO2/MWh] is the hourly marginal emission factor, and [€/tonne CO2] is the country‐specific social cost of carbon. Thus, the summation over and captures the total avoided CO2 emissions, weighted by the temporally varying marginal emission intensity and regionally differentiated damage valuation.
@@ A.1.8 Reduction in air pollutant (RAP)
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and transport.
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and optimized transport. We model this benefit as
where [kg/km] and [kg/km] are pollutant emission factors for EVs and ETs in country at time , reflecting local fleet composition and technology. and [km/year] are annual driving distances per EV and ET, and and are the respective vehicle counts. Thus, and give the pollutant reductions (kg) attributable to AI-enabled charging and scheduling.
The term [kg/MWh] denotes the hourly marginal emission factor for pollutant , while , , captures the RES output shifted and decarbonized through ODP optimization.
Finally, [€/kg] represents the country-specific marginal damage cost of pollutant , accounting for local air quality, health impacts, and income levels. The summation over and ensures that spatial and temporal heterogeneity in exposure and generation mix is reflected in the monetized benefit assessment, improving granularity and policy relevance relative to static average values.
| Name | Symbol | Unit | Value | Source |
| Baseline downtime cost | M€/yr | National statistics. | ||
| Self-efficiency improvement | % | Assumption (tested in sensitivity). | ||
| ODP downtime savings rate | % | Assumption. | ||
| Avg. electricity price | €/MWh | ENTSO-E [2025]. | ||
| Arbitrage energy share | % of energy | Assumption (proxy uplift). | ||
| RES capacity | MW | ENTSO-E [2025]. | ||
| RES full-load hours | h/yr | Eurostat [2025], European Environment Agency [2024]. | ||
| EV charging capacity | MW | dataset (Table 14); national stock data per Eurostat Eurostat [2025]. | ||
| ET charging capacity | MW | dataset (Table 14). | ||
| DR unit saving | €/MWh | National regulator tariff schedules. | ||
| Peak-shiftable share | % of energy | Assumption. | ||
| Avoided curtailment share | % of RES | Assumption. | ||
| Curtailment valuation | €/MWh | Assumption (opportunity cost proxy). | ||
| Fleet savings factor | % | Conservative assumption. | ||
| Per-EV annual saving | €/vehyr | IEA Global EV Outlook International Energy Agency [2024]. | ||
| Per-ET annual saving | €/vehyr | Rodríguez et al. [2023]. | ||
| Grid-stability improvement | % | Badanjak and Pandžić [2021]. | ||
| Fuel value (avoided) | €/MWh | ACER and CEER [2024]. | ||
| Grid-stability savings rate | % | Assumption. | ||
| Market price (stability) | €/MWh | ACER and CEER [2024]. | ||
| Value of lost load | €/MWh | CEPA study for ACER Cambridge Economic Policy Associates (2018) [CEPA]. | ||
| CO2 reduction factor | % | Assumption. | ||
| Emission factor (grid) | tCO2/MWh | European Environment Agency [2024]. | ||
| CO2 price | €/tCO2 | European Energy Exchange (2025) [EEX]. | ||
| Pollutant emission factor | kg/MWh | EEA and EMEP/CLRTAP [2019]. | ||
| Pollutant unit cost | €/kg | European Environment Agency [2020]. |
@@ A.3 OPEX breakdown
The figures in Table 12 represent the discounted present value (PV) of recurring annual costs OPEX over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The Annual average PV provides an annualized perspective of these costs, calculated from the total 10-year discounted OPEX. Costs specifically related to sustaining and evolving the AI engine are explicitly considered.
@@
To increase transparency in how AI-related expenditures are represented, we distinguish between one-time CAPEX (Table 11) and recurring OPEX (Table 12). CAPEX categories capture the upfront development and integration of AI modules, while OPEX categories cover recurring costs such as cloud hosting, AI inference and retraining, cybersecurity, and customer analytics. This breakdown clarifies that AI-specific costs are explicitly embedded in both cost categories, consistent with procurement benchmarks from AWS, Azure, Eurostat ICT reports, and IEA/ACER market integration studies. Table 13 summarizes the unit assumptions for AI model training, cloud infrastructure, and maintenance. These values are based on European and international benchmarks and were scaled using adoption and growth assumptions presented in Appendix A2. Together, they account for approximately 20% of total OPEX and 15% of total CAPEX.
| AI function / resource | Unit cost assumption | Source |
| AI model training (grid + transport) | €1-1.2 million per comprehensive model | [AWS, 2024, Azu, 2024] |
| AI inference (real-time predictions) | €0.05–0.07 per CPU-hour | [AWS, 2024, Azu, 2024] |
| Cloud storage and integration | €0.02–0.03 per GB-month | [Eurostat, 2023] |
| Model retraining and monitoring | €1.2–1.5M per model-year | [European Commission, 2023, International Energy Agency, 2023] |
| Cybersecurity and threat monitoring | €200–300k per country-year | [International Energy Agency, 2023] |
| API/market integration | €1.0–1.3M per country-year | [Agency for the Cooperation of Energy Regulators (2023), ACER] |
| Labor (AI/data engineers, ops staff) | €45–60k per FTE per year | [Eurostat, 2023] |
@@ B.3 Annual costs of the ODP (aggregated)
Note: Annual CAPEX and OPEX are discounted sums over the 2026–2035 period, aggregated across Austria, Hungary, and Slovenia. The Total row shows the ten‐year aggregate costs for each category. In addition, a one‐time upfront fixed cost of €89.6 million is incurred in 2026 to cover connectivity (5G/LPWAN), communication protocols and message queues, databases and data lakes, data integration middleware, market integration middleware, AI models for grid optimization, demand response tools, route optimization engines, weather-forecasting algorithms, software deployment and integration.